EVA Report – December 28th

Report: Crew 185 EVA #8 for SOL 10.

Propose of EVA:
-Collecting geological samples for soil density study. We will do multiple stops along Cow Dung road from the Hab to Gateway Lith. -We have done two stops along the way to collect samples, and it was a good teamwork!
-We stayed at the Lith for about 30 min for exploration and observing the panorama!
-Astro-CapCom communication. CapCom will drive the Astros to a specific geographical position (nearby the Hab) without the use of a GPS. EVA #8:
Start time: 1030am
End time: 1330pm->1430pm
Ilaria: Leader
Arno and John: Astros
David and Thibo are Comms.
Destination: Gateway to Lith (110.81 W – 38.45 N), see attachment. Vehicles: ATV (Arno), ATV (John), Rover Spirit (Ilaria)

Prepared by Ilaria Cinelli

EVA Report – December 26th

Propose of EVA: We did one three-hour EVA for three crewmembers. The purpose of this EVA was double:
– We wanted to test an innovative drone controller and a new GPS for analog spacesuits in the region of Pooh’s Corner (drone’s weight < 500g) (1h). Unfortunately, we didn’t manage to make these experience work in the field despite extensive efforts. We fixed them and we will try again tomorrow.
– We wanted to continue our experiment about control/astro communication. Arno and John looked for hidden objects with only radio indications and hid new objects to be found by Thibault and David.

EVA #6:
Start time: 1030am
End time: 0130pm

Arno Passeron: Leader
Thibault Paris: Astro 1

John Scezpaniak: Doctor Astro

Ilaria Cinelli and David Murray on Comms

Destination: Pooh’s Corner (E519000, N4251000) for part one then 200m south from the Hab (E518200, N4250500) for part two .



EVA #5 Report – December 24th

Report: Crew 185 EVA #5 for SOL 6.

Propose of EVA: We did a three-hour EVA this morning and we managed to do a lot of thing:

– We tested X-1 in a rocky environment
– We set up an experiment about soil erosion in an analog environment
– We continued our experiment about control/astro interaction
– We continued our experiment about the maximum weight that you can put on an ATV depending on the local soil density

EVA #5:

Start time: 1030am

End time: 0150pm

Ilaria Cinelli: Leader

David Murray: Astro 1

Thibault Paris: Astro 2

Arno Passeron and John Scezpaniak on Comms

Destination: Robert’s Rock Garden then Hab Area

Vehicles: 2

ATV: 350
Rover: Curiosity

Prepared by T. Paris

EVA Report – December 23rd

Report: Crew 185 EVA #4 for SOL 5.
Propose of EVA: Collecting geological samples for mobility testing. Then, walking with the EVA suits and discover the terrain around our position.

EVA Report:
(i) The first 40 min were dedicated to the engineering check during which we have filled the static tank with the left unfrozen water of the tanks on the trailer. Then, using a pan, we collected other (frozen) water that can be used for the Green Hab.
(ii) Arrived at the White Moon, we did a short walk around. Because strangers were near the vehicles, we went back, and we collected a few geological samples.
(iii) Then, we stopped about 1 km from the Hab to locate recognition signs to be used for the following two EVAs.
(iv) Back at the Hab, we refilled the ATV 350 and plugged the rover.

EVA #5:
Start time: 1030am
End time: 1230pm -> finished at 2pm

Ilaria Cinelli: Leader
John : Astro 1
Arno: Astro 2
Destination: White Moon (110.8W,38.44N).
Vehicles: 1 ATV (Ilaria), 1 Rover (John and Arno)
CapCom(s): David and Thibo

Written by Commander Ilaria Cinelli

EVA Report – November 26th

EVA #4 Report

The preparation of the 3D scanning device took longer than expected this morning, so we weren’t able to begin depressurization until 1:00 PM. Jordan, Ashton, Toruwa, and I hiked up to the top of Hab Ridge, searching for rocks that had enough detail. We surveyed the area for suitable specimens, but our scans kept running into problems. Since the regolith in the background for the loose boulders have a fairly uniform pattern and color, the scanner had trouble determining what it was supposed to scan. We’ll have to bring our own background with a randomized pattern.

Another issue that we ran into is that the practicality of using the 3D scanner on the field. Since it has to be attached to a laptop via a USB chord, we’d need one person to carry it, and another person to follow close behind while monitoring the status of the 3D scan.

The time that we were out took a surprisingly long time, by the time we descended the hill, it was about 3:00 PM. On the way down, Toruwa lost his footing and fell on his posterior. The fall seemed to deactivate the airflow fan in his suit, meaning we had to get him into the hab fast. By the time we got into the airlock, the inside of his helmet was already fogging up a lot. Since there were four of us, I went into the airlock with Toruwa, while Jordan and Ashton stayed outside, since they needed to further test the 3D scan equipment near the hab.

Due to the circumstances, the Commander ordered an emergency recompression, which took two minutes. Once the recompression was complete, we quickly opened the inner airlock door, getting his helmet off as quickly as we could. He was sweating profusely, and quite exhausted from the ordeal.

Jordan and Ashton’s activities lasted for another 20 minutes, reentering the hab at around 3:25 PM.

We succeeded in achieving what we needed to do on this EVA, but it’s really important to remember how risky it is to go on an EVA, because all it takes is one misstep, and the life support systems could go out.

This is all I have to report.

Thanks Mission Support!

Prepared by Brandon Ferguson

EVA report – November 25th

EVA #3 Report of Sol 3

EVA briefing took longer than expected. The fans on Space Suits #1 and #4 didn’t seem to be working properly. The LED lights for the attached chargers are currently displaying a red light. This will be touched on further in the engineering report. The stated purpose of our EVA was to test out the 3-D scanner on the field. Since it’s not operating properly at this time, we decided to bring two specimens back to the hab, so that they can be scanned here at a later time.

We finished decompression at 12:40 PM. We received a new shipment of gasoline for ATVs before our EVA. Ashton refueled all the ATVs from our activities from Sol 2. He completed refueling by about 12:50 PM. Toruwa and I took the rover, and Dean and Ashton followed in the ATVs. The Journey to the Stream Connector stop area took approximately 5 minutes

Once we parked, we walked about 50 meters to our first specimen, a softball-sized piece of green sandstone. We then found another suitable specimen, about the size of a small firewood log. We photographed and marked each site we extracted from, with plans to return these samples to there original locations at a later time. We returned to our Rovers by 1:20 PM, and began our recompression by 1:30 PM at the Hab.

We need to be clearer on the decompression / recompression time, as the inner airlock door was opened one minute earlier than it should have. I would recommend that the “official” decompression time be tracked by the occupants of the hab, as soon as the go-ahead is given by the EVA party.

Conditions outside were much warmer than expected. It’s a good thing our walking activities lasted for only about 40 minutes.

Communications between the members of the EVA party were concise and effective. Our Commander gave us phonetic alphabet designations based on our first name; for example Dean would be “Delta” and Ashton “Alpha”.

The EVA went smoothly and we accomplished all that we set out to do.

This is all I have to report.
Have a good night mission support!

Prepared by Brandon Ferguson
EVA Officer

EVA Report – November 24th

EVA #2 Report of Sol 2

We started our EVA at 11:50 AM. Jordan, Ashton, Toluwa, and I, each took an ATV, carrying cameras and other supplies in backpacks. We first stopped near the Stream Turnout Rd, and walked about 200 meters to an area with a large amount of dinosaur bones and petrified wood specimens. We spent about an hour in this area before heading north.

We arrived at the Lith Canyon turnout by about 12:20 PM. From here, we walked into the Lith Canyon area, careful not to step in the cactus gardens along the way.

We spent 2 hours here, exploring and photographing this area. There is a feature that is labeled on the map, called “The Pillar”, where the Copernicus Hwy North used to extend to. Although there are many features in this area that could fit this description, we found one particularly prominent column that could be “The Pillar”. We’ll have to investigate this further.

We returned to our ATVs by around 3:00 PM, getting back to the hab by 3:25 PM.

Have a great night Mission Support!

Prepared by Brandon Ferguson

EVA Report – November 23rd

EVA #1 Report

Today we conducted an out-of-sim EVA to the Dinosaur Quarry area, starting at approximately 3:00 PM. Ashton, Toluwa, Jordan, and I went on this EVA, each of us on our own ATV. The main purpose of today’s EVA was to drive the Blue 300 ATV longer than 5 km, so that its battery could recharge.

Before we set out, we did our ATV / Rover training, which lasted for about 20 minutes. Then we had to jumpstart the Blue 300 ATV, which had a low battery. By about 3:00 PM, we had the blue ATV running and ready to go. It took approximately 15-20 minutes for us to get there.

Once we arrived, I had a bit of a scare trying to restart the Blue ATV. It wouldn’t start at all! Then I realized that it was still in High. So I switched it to neutral and it started without any problem.

We spent over an hour in this area, following the canyon from the quarry site until we ended up in Lith Canyon, at the site where my EVA for crew 182 was conducted. We passed by one of the sites where I previously took regolith samples from. Just around the corner from this site was a large pile of petrified wood. Amazingly, these specimens look like fresh kindling wood, ready for the fireplace.

By around 4:30 PM, we made our way back to the ATVs, drove back to the hab, and returned by approximately 4:50 PM.

Have a great night Mission Support!

Prepared by Brandon Ferguson

EVA Officer

EVA Report – November 17th

EVA #10 Report

Today’s EVA began at 10:00 AM, involving Camila, Julio, and I. The Survivalman crew was filming our EVA, including the planning process, communications tests, putting on the spacesuits, and our decompression. The goal of our EVA was to go to Lith Canyon, collect three regolith samples, and return to the Hab. Each of the EVA crew boarded an ATV, with Les Stroud on his own ATV, and his camera crew on Deimos. I had Camila lead the way out so we could keep our ATV caravan nice and tight. The trip there took about an hour, because they needed footage of us along several different stretches of Cow Dung Rd. We would hold our position as they moved ahead, then we would move forward once they signaled that they were ready.

Once we arrived at the end of the Lith Canyon road, we proceeded on foot for about a kilometer westward. We had to be careful with our route finding as this was an area that none of us had been to before. We made sure to photograph the landscape at certain points so it would be easier to find our location later. Once we descended into the canyon, we made our way to an exposure with red and white regolith layers, taking one sample from each type. Once finished we walked further down the canyon where there were several eolian rock formations. There we collected our last regolith sample from a white exposure, turned around, and headed back the way we came. We spent about 2 hours walking in the Lith Canyon area, returning to the ATVs by around 1:20 PM. They didn’t need any more footage for the trip back, so we made our way back to the Hab by around 1:50 PM.

Today’s EVA was abnormally warm, especially all the walking we had to do. This was the longest EVA that I’ve had to do on foot, and was very challenging. I’d estimate that we walked a total of about 2 km over uneven terrain. It’s amazing how heavy and uncomfortable the suits felt on the first EVA when we just walked around the Hab. Now I feel a lot better in the EVA suit, even when covering much longer distances. Knowing this, I’m glad we didn’t plan any really long EVAs during our first few days of Sim, as it does take some time to get used to those suits. I’d definitely recommend that crew members take their time getting used to the suits, as their bulkiness can be surprising.

For our EVA, we managed to obtain the regolith samples we needed and the Survivorman crew was able to record our EVA operations.

This is all I have to report for today’s EVA.

Have a great night Mission Support!


Prepared by Brandon Ferguson

EVA Officer

EVA Report – November 16th

EVA #9 Report
Today’s EVA would’ve started at 11 AM and ended at 2 PM. Our goal would’ve been to go to Lith Canyon, collect samples from the Salt Wash Member, and return to the Hab. Our EVA for today did not take place, since we needed additional time to clean up the Hab before Survivorman arrives. Also we needed to figure out our transfer arrangements so that we’ll be able to get to Salt Lake City in time to catch their flight. EVA request coming soon!
Prepared by Brandon Ferguson
EVA Officer